a train changes its speed uniformly from 60 mph to 30 mph in a distance of 1,500 ft. what is its acceleration


1
Expert's answer
2021-03-14T19:20:18-0400

For the motion with constant acceleration the distance is

"s(t) = v_0t + \dfrac{at^2}{2}."

The velocity may be written in form "v(t) = v_0 + at," so "t = \dfrac{v(t)-v_0}{a}" .

Therefore, "s = \dfrac{v_0(v(t)-v_0)}{a} + \dfrac{(v(t)-v_0)^2}{2a} = \dfrac{v(t)^2-v_0^2}{2a}"

So "a= \dfrac{v(t)^2-v_0^2}{2s} = \dfrac{(30\cdot1.4667)^2-(60\cdot1.4667)^2 }{2\cdot1500} = -1.94\,\mathrm{ft/s^2}"